OpenStudy (anonymous):

Let c be a constant, and let f(x) be the following function: (x^2+6x-16 (3x^2-6x+2 (fraction) ,x=/=2 f(x)=( (7x-5 , x=2 Find all values of c (if any exist) which make f(x) continuous at x=2.

OpenStudy (e.mccormick):

OK.... that is just hard to read. Lets see what we can do with that first.

OpenStudy (anonymous):

Oh sorry I forgot to mention I already solved the fraction part. I came out with 2.

OpenStudy (anonymous):

I'm just having some trouble with the 7c-5 part. How can I take the limit of 7c-5 when I'm only given x...?

OpenStudy (e.mccormick):

You are looking to see where it makes the function continuous.

OpenStudy (anonymous):

Yea...

OpenStudy (e.mccormick):

is the \(3x^2-6x+2\) over \(3x^2-6x+2\) or? Just trying to get a grasp on the problem first.

OpenStudy (anonymous):

Yea sorry.

OpenStudy (e.mccormick):

And all that is at \(x\ne 2\)

OpenStudy (anonymous):

Wait wait... it's... (x^2+6x-16)/(3x^2-7x+2) at x=/=2

OpenStudy (anonymous):

Is that clearer?

OpenStudy (e.mccormick):

\(f(x)\begin{cases} \dfrac{x^2+6x-16}{3x^2-7x+2} & \text{ if } x\ne 2 \\ 7x-5 & \text{ if } x= 2 \end{cases}\)

OpenStudy (anonymous):

Yea. That's it. I wasn't sure how to do all the symbols and stuff.

OpenStudy (anonymous):

No. The bottom is 7c-5

OpenStudy (e.mccormick):

Right click that and look at show math as tex.

OpenStudy (e.mccormick):

OK. Not much of a difference.

OpenStudy (e.mccormick):

Have you taken the limit of: \(\dfrac{x^2+6x-16}{3x^2-7x+2}\)

OpenStudy (anonymous):

Yea and I got 2...

OpenStudy (anonymous):

But if x=/=2 did I do something wrong?

OpenStudy (e.mccormick):

well, it is: \(\lim\limits_{x\rightarrow2 }\dfrac{x^2+6x-16}{3x^2-7x+2}\) that you need to look at. That is because your goal is "Find all values of c (if any exist) which make f(x) continuous at x=2." For it to be continuous, there must NOT be a gap ar x=2.

OpenStudy (anonymous):

If you just plug in you get 0/0. So I factored to get ((x+8)(x-2))/((3x-1)(x-2)), then cancelled the two x-2's and got (x+8)/(3x-1), then when I plugged 2 in I got 10/5, or 2. Is that right?

OpenStudy (e.mccormick):

\(\lim\limits_{x\rightarrow2 }\dfrac{x^2+6x-16}{3x^2-7x+2}\) \(\lim\limits_{x\rightarrow2 }\dfrac{(x+8)\cancel{(x-2)}}{\cancel{(x-2)}(3x-1)}\) \(\lim\limits_{x\rightarrow2 }\dfrac{(x+8)}{(3x-1)}\)

OpenStudy (anonymous):

And then you plug in 2 for all the x's... right?

OpenStudy (e.mccormick):

Yep. And that gets you, as you said, 2. So, (2,2) is a hole, or undefined. And you want to fill it. Now, what does that mean the second equation must equal at x=2?

OpenStudy (anonymous):

c=2....? The c throws me off...

OpenStudy (anonymous):

Or... oh wait. 7c-5=2? So then solve for c?

OpenStudy (anonymous):

7c=7, c=1?

OpenStudy (e.mccormick):

Yah.

OpenStudy (anonymous):

So then what would the answer be?

OpenStudy (anonymous):

Is it (2,1)....? Or...

OpenStudy (e.mccormick):

c=1, as you said.

OpenStudy (anonymous):

That's the final answer?

OpenStudy (anonymous):

Oh I get it. Find all values of c. So c=1 is the answer...

OpenStudy (e.mccormick):

Yep. 1. See, you are looking for the value of c that makes f(x)=2 a valid equation. That is it.

OpenStudy (anonymous):

Thanks so much. I've been stuck on this for a while. I need to start taking better notes. :P

OpenStudy (e.mccormick):

np. Limits are an odd concept. They are not where you are, but where you are going. Wait until you learn that \(\frac{1}{3}=\infty\) in one respect.

OpenStudy (anonymous):

Lol I can't wait. :P Thanks again.

OpenStudy (e.mccormick):

It has to do with limits. If you take .3 + .03 + .003 + ... and so on, you get .3 that repeats until infinity. And that is 1/3rd. You take a limit of a sum function to prove this.

OpenStudy (anonymous):

Oh. o.o That's weird. I don't like math lol.

OpenStudy (anonymous):

I'm gonna close this question and open a new one, lol. Got a few more problems I can't finish.

OpenStudy (e.mccormick):

Kk.